como se hace esta operacion 1/a+1/b
Answer by Theo(13342)


multiply 1/a by b/b to get b/(ab)

multiply 1/b by a/a to get a/(ab)

now that the denominators are the same, you can combine the numerators under the same denominator.

you get:

b/(ab) + a/(ab) = (b+a) / (ab) = (a+b) / (ab)

your fraction becomes (a+b) / (ab)